# Peng Xia1, Ze Chen, Juanxi Tian, Yangrui Gong, Ruibo Hou2, Yue Xu Zhenbang Wu2, Zhiyuan Fan, Yiyang Zhou1, Kangyu Zhu3, Wenhao Zheng1 Zhaoyang Wang1, Xiao Wang4, Xuchao Zhang5, Chetan Bansal5

### v1

@@ Abstract
Artificial intelligence has significantly impacted medical applications, particularly with the advent of Medical Large Vision Language Models (Med-LVLMs), sparking optimism for the future of automated and personalized healthcare. However, the trustworthiness of Med-LVLMs remains unverified, posing significant risks for future model deployment. In this paper, we introduce CARES and aim to Comprehensively evAluate the tRustworthinESs of Med-LVLMs across the medical domain. We assess the trustworthiness of Med-LVLMs across five dimensions, including trustfulness, fairness, safety, privacy, and robustness. CARES comprises about 41K question-answer pairs in both closed and open-ended formats, covering 16 medical image modalities and 27 anatomical regions. Our analysis reveals that the models consistently exhibit concerns regarding trustworthiness, often displaying factual inaccuracies and failing to maintain fairness across different demographic groups. Furthermore, they are vulnerable to attacks and demonstrate a lack of privacy awareness. We publicly release our benchmark and code in https://github.com/richard-peng-xia/CARES.
@@ 1 Introduction
Artificial Intelligence (AI) has demonstrated its potential in revolutionizing medical applications, such as disease identification, treatment planning, and drug recommendation [43, 50, 58, 21, 10, 9, 47]. In particular, the recent emergence of Medical Large Vision Language Models (Med-LVLMs) has significantly enhanced the quality and accuracy of medical diagnoses [25, 37, 44, 13, 48], enabling more personalized and effective healthcare solutions. While Med-LVLMs have shown promising performance, existing models introduce several reliability issues [41, 51, 27, 54], including generating non-factual medical diagnoses, overconfidence in generated diagnoses, privacy breaches, health disparities, etc. The deployment of unreliable models can lead to severe adverse consequences [49, 33]. For instance, a model mistakenly identifying a benign tumor as malignant could lead to unnecessary invasive procedures and significant emotional distress for patients. Therefore, understanding and evaluating the trustworthiness of Med-LVLMs is paramount in medical applications.
@@ 2 CARES Datasets
Construction of QA Pairs. We explore the processes of constructing QA pairs from both closed-ended and open-ended questions. Firstly, we delve into closed-ended questions. For closed-ended yes/no questions, we utilize the OL3I [61] and IU-Xray [6] datasets, converting their questions along with corresponding labels or reports into yes/no formats. For example, the question "Can ischemic heart disease be detected in this image?" is transformed accordingly. For closed-ended multi-choice questions, the multi-class classification dataset HAM10000 [45] is converted into QA pairs with multiple options. For example, in the HAM10000 [45] dataset, for lesion types, we can design the following QA pair: Question: What specific type of pigmented skin lesion is depicted in this dermatoscopic image? The candidate options are:[A:melanocytic nevi, B:dermatofibroma, C:melanoma, D:basal cell carcinoma]; Answer: A:melanocytic nevi. To increase the diversity of question formats and ensure the stability of testing performance, we design 10-30 question templates for multi-choice question type (see detailed templates in Appendix C). Furthermore, to enrich the dataset with diverse modalities and anatomical regions, a comprehensive multi-choice VQA dataset, OmniMedVQA [15] is also collected. For open-ended questions, CARES features a series of open-ended questions derived from vision-language datasets, namely MIMIC-CXR [19], Harvard-FairVLMed [35], and PMC-OA [28]. Specifically, medical reports or descriptions are transformed into a series of open-ended QA pairs by GPT-4 [39] (see details in Appendix C).
@@ 3.3 Safety Evaluation and Results
Similar to LVLMs [46], Med-LVLMs also present safety concerns, which include several aspects such as jailbreaking, over-cautious behavior, and toxicity. Addressing these issues is paramount to ensuring the safe deployment of Med-LVLMs.
@@ 3.5 Robustness Evaluation and Results
Table 5: Abstention rate (Abs) and accuracy (ACC) (%) tested on noisy data.
@@
ACC Abs ACC Abs
LLaVA-Med 57.28 $\downarrow$9.33 6.05 28.49 $\downarrow$6.21 7.31
Med-Flamingo 23.29 $\downarrow$3.45 0 51.70 $\downarrow$10.20 0
MedVInT 64.38 $\downarrow$8.96 0 51.47 $\downarrow$10.43 0
RadFM 25.29 $\downarrow$1.38 0.02 19.04 $\downarrow$1.46 0.01
@@ 4 Related Work
Medical Large Vision Language Models. LVLMs have demonstrated remarkable performance in natural images [39, 67, 31, 1], which has facilitated their application in the medical domain. Recent advancements have witnessed the emergence of Med-LVLMs such as LLaVA-Med [25] and Med-Flamingo [37]. They are built upon the foundation of open-source general LVLMs, subsequently fine-tuned using biomedical instruction data across various medical modalities. Additionally, several Med-LVLMs tailored to specific medical modalities have been developed, such as XrayGPT [44] (radiology), PathChat [34] (pathology), and OphGLM [8] (ophthalmology). These models hold immense potential to positively impact the healthcare field, e.g., by providing reliable clinical recommendations to doctors. As LVLMs are deployed in increasingly diverse fields, concerns regarding their trustworthiness are also growing [42, 49], particularly in the medical field. Unreliable models may induce hallucinations and results in inconsistencies between image-textual facts [26] or may result in unfair treatment based on gender, race, or other factors [35]. Hence, proposing a comprehensive trustworthiness benchmark for Med-LVLMs is both imperative and pressing.
Trustworthiness in LVLMs. In LVLMs, existing evaluations of trustworthiness primarily focus on specific dimensions [33, 57], such as trustfulness  [26, 7, 24, 57, 59, 5, 53, 64] or safety [46, 40]. Specifically, for trustfulness, LVLMs may suffer from hallucinations that conflict with facts [65, 66, 52, 4]. Previous methods evaluate LVLM hallucinations for VQA [26, 7, 11] and captioning [26, 5, 53, 64], with models exhibiting significant hallucinations. For safety, attack and jailbreak strategies are leveraged to induce erroneous responses [46]. Similarly, Med-LVLMs inherit these issues of trustfulness and safety, as indicated by single-dimension evaluations [41, 27]. Unlike these studies that mainly focus on a specific dimension, we are the first to conduct a holistic evaluation of trustworthiness in Med-LVLMs, including trustfulness, fairness, safety, privacy, and robustness.
@@ Appendix C Construction Process of QA Pairs
Summary. After constructing QA pairs, the data utilized in CARES is summarized as shown in Table 13. These statistics reveal that CARES includes 18K images and 41K question-answer pairs, encompassing a variety of question types and covering 16 medical image modalities and 27 human anatomical regions. Moreover, to better present the diversity of medical image modalities and anatomical regions, we illustrate the images with the corresponding QA items in Figure 7.

### v3

# Peng Xia1,2, Ze Chen2, Juanxi Tian2, Yangrui Gong2, Ruibo Hou7, Yue Xu2 Zhenbang Wu7, Zhiyuan Fan9, Yiyang Zhou1, Kangyu Zhu3, Wenhao Zheng1 Zhaoyang Wang1, Xiao Wang4, Xuchao Zhang5, Chetan Bansal5

@@ Abstract
Artificial intelligence has significantly impacted medical applications, particularly with the advent of Medical Large Vision Language Models (Med-LVLMs), sparking optimism for the future of automated and personalized healthcare. However, the trustworthiness of Med-LVLMs remains unverified, posing significant risks for future model deployment. In this paper, we introduce CARES and aim to Comprehensively evAluate the tRustworthinESs of Med-LVLMs across the medical domain. We assess the trustworthiness of Med-LVLMs across five dimensions, including trustfulness, fairness, safety, privacy, and robustness. CARES comprises about 41K question-answer pairs in both closed and open-ended formats, covering 16 medical image modalities and 27 anatomical regions. Our analysis reveals that the models consistently exhibit concerns regarding trustworthiness, often displaying factual inaccuracies and failing to maintain fairness across different demographic groups. Furthermore, they are vulnerable to attacks and demonstrate a lack of privacy awareness. We publicly release our benchmark and code in https://cares-ai.github.io/.
@@ 1 Introduction
Artificial Intelligence (AI) has demonstrated its potential in revolutionizing medical applications, such as disease identification, treatment planning, and drug recommendation [60, 67, 84, 29, 14, 13, 64, 77, 19, 20, 34]. In particular, the recent emergence of Medical Large Vision Language Models (Med-LVLMs) has significantly enhanced the quality and accuracy of medical diagnoses [33, 47, 61, 17, 65], enabling more personalized and effective healthcare solutions. While Med-LVLMs have shown promising performance, existing models introduce several reliability issues [51, 68, 37, 72], including generating non-factual medical diagnoses, overconfidence in generated diagnoses, privacy breaches, health disparities, etc. The deployment of unreliable models can lead to severe adverse consequences [66, 43]. For instance, a model mistakenly identifying a benign tumor as malignant could lead to unnecessary invasive procedures and significant emotional distress for patients. Therefore, understanding and evaluating the trustworthiness of Med-LVLMs is paramount in medical applications.
@@ 2 CARES Datasets
Construction of QA Pairs. We explore the processes of constructing QA pairs from both closed-ended and open-ended questions. Firstly, we delve into closed-ended questions. For closed-ended yes/no questions, we utilize the OL3I [90] and IU-Xray [10] datasets, converting their questions along with corresponding labels or reports into yes/no formats. For example, the question "Can ischemic heart disease be detected in this image?" is transformed accordingly. For closed-ended multi-choice questions, the multi-class classification dataset HAM10000 [62] is converted into QA pairs with multiple options. For example, in the HAM10000 dataset, for lesion types, we can design the following QA pair: Question: What specific type of pigmented skin lesion is depicted in this dermatoscopic image? The candidate options are:[A:melanocytic nevi, B:dermatofibroma, C:melanoma, D:basal cell carcinoma]; Answer: A:melanocytic nevi. To increase the diversity of question formats and ensure the stability of testing performance, we design 10-30 question templates for multi-choice question type (see detailed templates in Appendix C). Furthermore, to enrich the dataset with diverse modalities and anatomical regions, a comprehensive multi-choice VQA dataset, OmniMedVQA [21] is also collected. For open-ended questions, CARES features a series of open-ended questions derived from vision-language datasets, namely MIMIC-CXR [27], Harvard-FairVLMed [45], and PMC-OA [38]. Specifically, medical reports or descriptions are transformed into a series of open-ended QA pairs by GPT-4 [49] (see details in Appendix C).
@@ 3.3 Safety Evaluation and Results
Similar to LLMs [83] and LVLMs [63], Med-LVLMs also present safety concerns, which include several aspects such as jailbreaking, over-cautious behavior, and toxicity. Addressing these issues is paramount to ensuring the safe deployment of Med-LVLMs.
@@ 3.5 Robustness Evaluation and Results
Table 5: Abstention rate (Abs), accuracy (ACC) and AUROC (%) tested on noisy data.
@@
ACC AUROC Abs ACC AUROC Abs
LLaVA-Med 57.28 $\downarrow$9.33 59.83 $\downarrow$6.87 6.05 28.49 $\downarrow$6.21 52.30 $\downarrow$4.43 7.31
Med-Flamingo 23.29 $\downarrow$3.45 52.14 $\downarrow$3.22 0 51.70 $\downarrow$10.20 59.47 $\downarrow$8.20 0
MedVInT 64.38 $\downarrow$8.96 65.82 $\downarrow$8.42 0 51.47 $\downarrow$10.43 58.82 $\downarrow$9.38 0
RadFM 25.29 $\downarrow$1.38 53.69 $\downarrow$1.62 0.02 19.04 $\downarrow$1.46 50.56 $\downarrow$0.69 0.01
@@ 4 Related Work
Medical Large Vision Language Models. LVLMs have demonstrated remarkable performance in natural images [49, 97, 41, 1, 58, 94, 75, 78, 69, 52, 24, 25, 4, 5, 71, 86], which has facilitated their application in the medical domain. Recent advancements have witnessed the emergence of Med-LVLMs such as LLaVA-Med [33] and Med-Flamingo [47]. They are built upon the foundation of open-source general LVLMs, subsequently fine-tuned using biomedical instruction data across various medical modalities. Additionally, several Med-LVLMs tailored to specific medical modalities have been developed, such as XrayGPT [61] (radiology), PathChat [44] (pathology), and OphGLM [12] (ophthalmology). These models hold immense potential to positively impact the healthcare field, e.g., by providing reliable clinical recommendations to doctors. As LVLMs are deployed in increasingly diverse fields, concerns regarding their trustworthiness are also growing [59, 66, 80, 79], particularly in the medical field. Unreliable models may induce hallucinations and results in inconsistencies between image-textual facts [36] or may result in unfair treatment based on gender, race, or other factors [45]. Hence, proposing a comprehensive trustworthiness benchmark for Med-LVLMs is both imperative and pressing.
Trustworthiness in LVLMs. In LVLMs, existing evaluations of trustworthiness primarily focus on specific dimensions [43, 82], such as trustfulness  [36, 11, 32, 82, 85, 9, 70, 56, 55, 57, 53, 54, 81] or safety [63, 50, 7, 6]. Specifically, for trustfulness, LVLMs may suffer from hallucinations that conflict with facts [95, 96, 69, 8, 87, 91, 88, 35, 76]. Previous methods evaluate LVLM hallucinations for VQA [36, 11, 15] and captioning [36, 9, 70, 94], with models exhibiting significant hallucinations. For safety, attack and jailbreak strategies are leveraged to induce erroneous responses [63]. Similarly, Med-LVLMs inherit these issues of trustfulness and safety, as indicated by single-dimension evaluations [51, 37]. Unlike these studies that mainly focus on a specific dimension, we are the first to conduct a holistic evaluation of trustworthiness in Med-LVLMs, including trustfulness, fairness, safety, privacy, and robustness.
@@ Acknowledgement
Appendix
WARNING: The Appendix contains model outputs that may be considered offensive.
@@ Appendix C Construction Process of QA Pairs
Summary. After constructing QA pairs, the data utilized in CARES is summarized as shown in Table 13. These statistics reveal that CARES includes 18K images and 41K question-answer pairs, encompassing a variety of question types and covering 16 medical image modalities and 27 human anatomical regions. Moreover, to better present the diversity of medical image modalities and anatomical regions, we illustrate the images with the corresponding QA items in Figure 7. Moreover, we illustrate some cases in Figure 8 to provide a more intuitive understanding of trustworthiness issues in Med-LVLMs.